OpenStudy (goformit100):

If Jean’s weekly income doubled she would be making $120 a week more than Barbara. Jean’s weekly income is $80 more than half of Betty’s. Betty makes $200 a week. How much does Barbara make? a. $180 c. $240 b. $200 d. $360

OpenStudy (goformit100):

@Noemi95 @TessaMarie08181997

OpenStudy (anonymous):

Let J be jean's income, A be barbara's, and E be betty's 2J = 120+A J = 80 + E/2 E=200 ------------ J = 80+200/2 J = 80+100 J = 180 ------------ 2(180) = 120+A 360 = 120+A 360-120 = A 240 = A. ------------ So the answer is (c)

OpenStudy (anonymous):

yeah c